OpenStudy (linn99123):

need help fast will fan and give medal The lengths of the legs of a right triangle are 25 cm and 60 cm. What is the length of the hypotenuse? A. 9.2 centimeters B. 42.5 centimeters C. 65 centimeters D. 35 centimeters

OpenStudy (abmon98):

(25)^2+(60)^2=4225 take the square root of 4225=65 centimeters

OpenStudy (shamil98):

Use the Pythagorean theorem to solve for the hypotenuse. \[a^2 + b^2 = c^2\] Where a and b are the legs, and c represents the hypotenuse. \[25^2 + 60^2 = c^2\] Simplify the constants and solve for c.

OpenStudy (anonymous):

What sham said

OpenStudy (anonymous):